OpenStudy (anonymous):

20! / 16! = A. 24 B. about 16 C. 116,280

OpenStudy (asnaseer):

Do you know what factorial means, i.e. do you know what 20! translates to?

OpenStudy (anonymous):

Yes and i figured this problem would be easy for me, but the answer i got was quite far off of my answer choices

OpenStudy (asnaseer):

so how would you represent 20! ?

OpenStudy (asnaseer):

or take a simpler one, how would you represent, say, 4! ?

OpenStudy (anonymous):

20*19*18*17*16*15*14*13*12*11*10*9*8*7*6*5*4*3*2*1=14311182833920000

OpenStudy (asnaseer):

ok - I was only looking for the representation - not the worked out answer - but you have it correct

OpenStudy (asnaseer):

so, in your question we have:\[\frac{20!}{16!}=\frac{20*19*18*17*16*15*...*2*1}{16*15*...*2*1}\]

OpenStudy (asnaseer):

you should notice that a lot of the terms cancel out

OpenStudy (asnaseer):

this is because:\[20!=20*19*18*17*16*15*...*2*1=20*19*18*17*16!\]therefore the 16! cancels out in this fraction leaving you with just:\[20*19*18*17\]

OpenStudy (asnaseer):

do you follow?

OpenStudy (anonymous):

Yes

OpenStudy (asnaseer):

so you should be able to get the right answer now